How many liters of gaseous hydrogen bromide at 69°C and 0.889 atm will a chemist need if she wishes to prepare 3.50 L of 1.20 M hydrobromic acid
Answer #1

we know the equation

PV=nRT

V=nRT/P

here

P=0.889atm

T=273+69=342 kelvin

R=0.0821

n=3.50 x 1.20=4.2m

V=4.2 x 0.0821 x 342/0.889

V=117.92/0.889

V=132.65 L

ANS:132.65 liters of gaseous hydrogen bromide at 69°C and 0.889 atm will a chemist need if she wishes to prepare 3.50 L of 1.20 M hydrobromic acid